One roots of the equation x^3 - 2x^2 – 13x - 10 = 0 is 5. What is the sum of its other roots?
(a)-2 (b) -3 (c) 2 (d) 1 (e) None of these

Answer by stanbon(75887)

One root of the equation x^3 - 2x^2 – 13x - 10 = 0 is 5. What is the sum of its other roots?
----
Use synthetic division:
5)....1....-2....-13....-10
......1.....3.....2....|..0
Quotient: x^2 + 3x + 2
---
Solve x^2 + 3x + 2 = 0
Factor: (x+2)(x+1) = 0
x = -1 or x = -2
Sum = -3
==========================
Cheers,
Stan H.
========================
(a)-2 (b) -3 (c) 2 (d) 1 (e) None of these